# Astromeric nature of 119mAg unveiled through direct mass measurement

### v1

@@ VIII Acknowledgments
This work was supported in part by the National Science Foundation (NSF) Grant No. PHY-2310059, U.S. Department of Energy by Lawrence Livermore National Laboratory under Contract DE-AC52-07NA27344, the US Department of Energy, and supported from the Natural Sciences and Engineering Research Council of Canada: Grant SAPPJ-2018-00028, as well as by the U.S. Department of Energy, Office of Nuclear Physics, under Contract No. DE-AC02-06CH11357 (ANL). L.V. was supported by the National Science Foundation Graduate Research Fellowship under Grant No. DGE-1746045. This research used resources of ANL’s ATLAS facility, which is a DOE Office of Science User Facility. M.R.M. acknowledges support from the Laboratory Directed Research and Development program of LANL project number 20230052ER. G.W.M., M.R.M., and T.M.S. were supported by the US Department of Energy through LANL. LANL is operated by Triad National Security, LLC, for the National Nuclear Security Administration of U.S. Department of Energy (Contract No.326 89233218CNA000001).

### v3

@@ Abstract
Acknowledgments - This work was supported in part by the National Science Foundation (NSF) Grant No. PHY-2310059, U.S. Department of Energy by Lawrence Livermore National Laboratory under Contract DE-AC52-07NA27344, the US Department of Energy, and supported from the Natural Sciences and Engineering Research Council of Canada: Grant SAPPJ-2018-00028, as well as by the U.S. Department of Energy, Office of Nuclear Physics, under Contract No. DE-AC02-06CH11357 (ANL). L.V. was supported by the National Science Foundation Graduate Research Fellowship under Grant No. DGE-1746045. This research used resources of ANL’s ATLAS facility, which is a DOE Office of Science User Facility. M.R.M. acknowledges support from the Laboratory Directed Research and Development program of LANL project number 20230052ER. G.W.M., M.R.M., and T.M.S. were supported by the US Department of Energy through LANL. LANL is operated by Triad National Security, LLC, for the National Nuclear Security Administration of U.S. Department of Energy (Contract No. 89233218CNA000001).